OpenStudy (anonymous):

Find all solutions in interval (0,2pi) 2 sin^2 (x) + cos (x) = 1 Write in terms of pi PLEASE HELP!

OpenStudy (anonymous):

the answer has to be in radians (in terms of pi)

OpenStudy (anonymous):

hi guys! Please help! I will truly appreciate it!

OpenStudy (anonymous):

180 degrees

OpenStudy (anonymous):

? is that it?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

how? in terms of pi?

OpenStudy (anonymous):

2 sin^2 x - cos x = 1 2(1-cos^2 x) - cos x = 1 where 1=sin^2x + cos^2x 2 - 2cos ^2x - cos x = 1 expand and collect like terms -2cos^2x - cos x = -1 2cos^2 x + cos x = 1 multiply by negative one (both sides) 2 cos^2 x + cos x - 1 = 0 quadratic equation You can refer to 'PIANOs' excellent method below, which works with the radical, and always yields an exact answer. You could have also factored this one: (2u - 1)(u+1)= 0 could have been used where u = cosx cosx = 1/2 cosx= -1 So, the 'angles' that jive with that cosine ratio are: cos x = 1/2 cos^-1 (cosx) = cos^-1 (1/2) x=60 degrees, or 300 degrees cosx = -1 cos^-1(cosx) = cos^-1(-1) x = 180 degrees.

OpenStudy (anonymous):

so is the solution just 180 or is there more? and 180 in terms of radian/pi is just pi right?

OpenStudy (anonymous):

The question asked for 2 sin^2 x + cos x = 1, not 2 sin^2 x - cos x = 1

OpenStudy (anonymous):

yeah its + not subtraction!

OpenStudy (anonymous):

can u help me rich_m_c

OpenStudy (anonymous):

Check out this plot and try to tell me the solutions: http://www.wolframalpha.com/input/?i=2sin%5E2%28x%29%2Bcos%28x%29%3D1+from+0+to+2*pi

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i dont understsand it

OpenStudy (anonymous):

The curve is y=2 sin^2 (x) + cos (x) and where the red line crosses at the red points is y=1

OpenStudy (anonymous):

please i am so desperate!

OpenStudy (anonymous):

but how do i put the solutions in terms of pi?

OpenStudy (anonymous):

Pleasee!

OpenStudy (anonymous):

The graph was just an aid to the eye. To find the solution you have to follow a similar method to Drakula's